ELiGiO:

Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1) Enter your answer into the blank equation box.

snowflake0531:

You're offline, so I'll just do this Firstly, you should know that we want to find slope and y-intercept in slope intercept, adn then change it to standard form Slope intercept form: y=mx+b Standard form: ax + by=c And so, to first calculate the slope, we'll do

1 attachment
snowflake0531:

So we plug P(0, -4), Q(5, 1) in \[\frac{ 1-(-4) }{ 5-0 }\] we can simplify this to \[\frac{ 5 }{ 5 } = 1 So, slope is 1

snowflake0531:

y=mx+b m is slope, so y=x+b Now we have P(0, -4), Q(5, 1) We can choose one of the two to put into the equation y=x+b to find b, we can use the first for this one (0,-4) -4 = a+b b=-4

snowflake0531:

So we know that the slope intercept form of htis is y=x-4

snowflake0531:

So now we convert this to standard form: ax+by=c We can just but the x to the left, and get -x + y = -4

ELiGiO:

Thank you bro....i wish i was online when u did this but i thought nobody would even help me. But thanks bro